# Fast-Convergent Proximity Graphs for Approximate Nearest Neighbor Search

### v4

@@ 3.3.1. Space and Construction Time
    where the last equality used the assumption $d=O(1)$. Since the minimum distance between any two points in $P$ is $1$, each ball of radius $\frac{1}{3}$ contains at most one point (see Figure 3(a)). Thus, $|\text{Ring}_{i}|=O((\alpha\cdot\tau)^{d})$, meaning that $p$ has $O((\alpha\cdot\tau)^{d})$ out-neighbors in $\text{Ring}_{i}$.
    (a) Case 1
    (b) Case 2
    Figure 3. Out-degree analysis in two cases
@@
    Similarly, we can obtain $\mathit{\delta}(p,v)>\mathit{\delta}(u,v)\cdot\alpha+(\alpha+1)\tau$. According to the pruning rule defined by (1), at most one of $u$ and $v$ can be in the shortcut set $S$ of $p$. Hence, for any ball $B$ with radius $r_{i}$, at most one point in $\text{Ring}_{i}\cap B$ can be an out-neighbor of $p$. For instance, in Figure 3(b), when $v$ and $v^{\prime}$ are the our-neighbors of $p$, then edges $(p,u)$ and $(p,u^{\prime})$ will not exist in our PG.
@@ 3.3.2. Query Time
    Now, we prove that $\mathit{\delta}(p_{\ell-1},q)\geq 1/2$. If $\mathit{\delta}(v^{*},q)\geq 1/2$, then, since $v^{*}$ is the exact NN of $q$ in $P$, we have $\mathit{\delta}(p_{\ell-1},q)>\mathit{\delta}(q,v^{*})>1/2$. On the other hand, if $\mathit{\delta}(v^{*},q)<1/2$, by definition of the metric space and the fact that the minimum pairwise distance in $P$ is $1$, we obtain
@@
    When $\mathit{\delta}(q,v^{*})>\tau$. We say a proximity graph $G^{\prime}$ is $\alpha$-shortcut reachable if for every two vertices $p,z$ of $G^{\prime}$ such that $(p,z)$ is not in $G^{\prime}$, then, there exists an edge $(p,p^{\prime})$ in $G^{\prime}$ such that $\mathit{\delta}(p^{\prime},z)\leq\mathit{\delta}(p,z)/\alpha$. Indyk and Xu (ix23, Theorem 3.4) proved that given an $\alpha$-shortcut reachable PG, a greedy routing starting at any vertex can answer an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN query after visiting $O(\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ hop vertices. Our $\alpha$-CG $G$ is also $\alpha$-shortcut reachable:
@@ 4. Alpha-Convergent Neighborhood Graph with Efficient Construction
    Despite the superior asymptotic query performance of the $\alpha$-CG, its construction time is $\Omega(n^{2})$ in the worst case. In line with existing works such as NSG (fxwc19), Vamana (sds+19), and $\tau$-MNG (pcc+23), we propose a practical variant that approximates $\alpha$-CG. This section also explores strategies for adaptively pruning candidates (setting the parameter $\alpha$) and constructing the index efficiently.
@@ 4.1. The Overall Framework
    We employ a standard framework utilized in existing PG methods, such as NSG (fxwc19) and $\tau$-MNG (pcc+23), to generate a local-neighbor set (as opposed to $P\setminus{p}$) as the candidate set for each point $p\in P$. Then, we apply an adaptive pruning strategy (Section 4.2) to select up to $M$ points as the out-neighbors of $p$, where $M$ is a predefined maximum out-degree that ensures efficient graph storage and computation. We refer to resulting graph $G$ as the $\alpha$-convergent neighborhood graph ($\alpha$-CNG). This framework consists of the following phases, as summarized in Algorithm 3.
    The initial phase constructs an approximate $K$-NN graph $G_{0}$, where each data point is connected to its approximate $K$-NNs in $P$. Efficient implementations are available for constructing such graphs, e.g., (hd16; dcl11). We then identify the navigating node $s$ by performing a beam search, starting from a random point with the centroid of $P$ (the geometric mean of all points in $P$) as the query point. The returned NN is designated as $s$.
@@
    Subsequently, backward edges are added to the graph for each edge inserted in the previous phase, ensuring bidirectional connectivity. If any vertex $p$’s out-degree exceeds $M$, its out-neighbor set is pruned accordingly using our adaptive pruning method. Finally, we verify the graph connectivity by performing a depth-first search (DFS) starting from $s$. To ensure that all nodes are reachable from $s$, we add necessary edges for any nodes not included in the DFS tree, in line with the NSG method (fxwc19).
@@ 4.3. Efficient Graph Construction
    Total construction time. The construction time of $\alpha$-CNG is dominated by two parts: (i) approximate $K$-NN graph construction, and (ii) candidates pruning for all $p\in P$. We utilize the NN-decent algorithm (dcl11) to compute the approximate $K$-NN graph, whose empirical time complexity is sub-quadratic.
@@ 4.4. Discussions
    Although $\tau$ is a problem parameter, in our experiments, we set $\tau$ to a small value within the range $[0,10]$ (and in most cases, $\tau$ is smaller than 1) to avoid the large maximum out-degree. Both $\tau$-MNG (pcc+23) and our paper found that the optimal settings for $\tau$ may vary across different datasets. Hence, we apply a grid search on the test queries. In practical applications, historical queries can be leveraged to determine this parameter. Specifically, we first compare the results obtained with $\tau=0$ against those with $\tau\in\{10,\,1,\,0.1,\,0.01,\,0.001\}$ to identify a coarse range of $\tau$, and then fine-tune $\tau$ to locate the best value.
    Updates. This paper focuses on the development of static PGs that enhance both query accuracy and query time. Supporting updates is beyond the scope of this work. Nonetheless, concepts from existing research—such as periodic global rebuilding (pcc+23), lazy deletion with masking (ssks21), and in-place update strategies (xmb+25)—could be applied to our $\alpha$-CNG and may be considered for future work.
    5. Experiments
@@
    SIFT 128 1M 10K (jds10) Image
    CRAWL 300 1.98M 10K (mgbpj18) Text
    WIKI 384 1M 1K (mxbs16) Text
    MSONG 420 1M 200 (bewl11) Audio
    LAION-I2I 768 1M 10K (sbv+22) Image
    GIST 960 1M 1K (jds10) Image
    DEEP100M 96 100M 10K (bl16) Image
    BIGANN100M 128 100M 10K (swa+22) Image
    5.1. Experimental Setup
@@
    Datasets. We utilized eight real-world datasets, which are widely adopted in ANN search evaluation (pcc+23; ggxl25; cscz24; fwc21; fxwc19; my20; yxl+24). These datasets include six at the 1M scale and two at the 100M scale. The 1M-scale datasets span various application domains, including image (SIFT, LAION-I2I, GIST), audio (MSONG), and text (WIKI, and CRAWL). We also assessed the scalability of our methods using the 100M-scale datasets, including DEEP100M [^5] DEEP100M consists of the first 100 million vectors from the public DEEP1B(bl16) dataset. and BIGANN100M. Table 2 summarizes the key statistics, including the dimensionality (Dim.), the number of base points (# Base), the number of query points (# Query), data source, and data type.
    Competing Methods. Our first method, $\alpha$-CNG, employs an adaptive pruning strategy (Section 4.2). To evaluate the effectiveness of this strategy, we examined Fixed-$\alpha$-CNG, a variant that employs a global $\alpha$ for all data points. Since previous studies (lzs+19; wxyw21) have consistently shown that PG-based methods outperform non-PG methods such as LSH and IVF, we focus our comparisons on PG-based baselines. We compared our methods against four state-of-the-art PG algorithms: HNSW (my20), Vamana (sds+19), NSG (fxwc19), and $\tau$-MNG (pcc+23), selected for their robust performance. Three additional popular algorithms—NSSG (fwc21), DPG (lzs+19), and FANNG (hd16)—were excluded as $\tau$-MNG (pcc+23) outperforms their performance. All baseline implementations use publicly available source code.
    The construction parameters for all structures were selected based on official recommendations and empirical evaluations. HNSW used a configuration of $M$ = 32 and $ef_{C}$ = 500 for all datasets, while Vamana used the settings in (sds+19), with $M$ = 70, $L$ = 75, and $\alpha$ = 1.2.
@@ 5.2. Search Performance
    Recall vs. NDC. We first assessed the trade-off between (average) recall and (average) NDC on the six datasets at the 1M scale. Figure 6 reports the results obtained by varying the queue size $L$ of beam search. Our $\alpha$-CNG outperformed all baselines, achieving higher recall with fewer NDCs. The only exception is the SIFT dataset, for which all methods exhibited strong performance due to its low intrinsic dimensionality (pcc+23). Fixed-$\alpha$-CNG also surpassed the four baselines, except on MSONG, but was outperformed by $\alpha$-CNG.
@@
    Recall vs. # hops. We next evaluated the (average) number of hops by varying $L$. Figure 6 shows that $\alpha$-CNG substantially reduced the number of hops across all datasets, and Fixed-$\alpha$-CNG also outperformed the baselines on most datasets. Table 3 confirms that the reductions in the number of hops over the best-performing baseline are substantial: exceeding 45% on five out of the six datasets, with a maximum speedup of 2.88x. These gains arise from two key factors: $\alpha$-CNG approximates $\alpha$-CG for faster convergence, and our adaptive pruning strategy preserves more long-distance shortcut edges, effectively reducing the number of hops. The reduced hops of $\alpha$-CNG are advantageous for disk-based (sds+19) or distributed deployments of PGs, where hops correspond to I/O operations.
@@ 5.3. Effect of Parameters
    Varying $\tau$. Figure 9 illustrates the impact of the parameter $\tau$ on search performance, where the middle tested $\tau$ for each dataset is the optimal $\tau$ identified. We observe a trend similar to that reported for $\tau$-MNG (pcc+23): as $\tau$ increases from zero, search performance initially improves before deteriorating with further increases in $\tau$. Recall that the search time is proportional to the total out-degrees of the hop vertices. A moderate increase in $\tau$ allows connecting more candidates and enhances the graph connectivity, thus reducing the number of hops. However, if $\tau$ increases excessively, the out-degree may become too large, as the search must compute distances for more neighbors at each step. Additionally, the increased out-degrees may exceed the limit $M$ for some nodes, resulting in the replacement of long-distance shortcut edges with less useful ones. Overall, a small positive $\tau$ enhances graph connectivity and improves search performance, while excessively large values can lead to over-inclusion and reduced pruning effectiveness.
@@ 5.4. Construction Performance
    Table 4. Index time (secs)
    Method SIFT CRAWL WIKI MSONG LAION-I2I GIST
    HNSW 42 159 91 145 116 206
    Vamana 9 43 19 38 35 67
    NSG 43 174 97 223 93 200
    $\tau$-MNG 43 182 98 242 98 217
    Fixed-$\alpha$-CNG 42 206 88 215 99 207
    $\alpha$-CNG 53 221 119 227 110 228
    Table 5. Index size (MB)
    Method SIFT CRAWL WIKI MSONG LAION-I2I GIST
    HNSW 256 509 256 254 256 256
    Vamana 163 409 114 151 134 118
    NSG 115 183 113 97 83 90
    $\tau$-MNG 113 329 124 117 127 131
    Fixed-$\alpha$-CNG 158 512 179 203 186 162
    $\alpha$-CNG 188 563 228 225 246 227
@@ 5.5. Effectiveness of Our Edge Pruning Rule
    Our two methods, NSG, and $\tau$-MNG, share a framework that generates candidate sets from approximate $K$-NN graphs. Section 5.2 shows the efficacy of our pruning rule (Definition 2) within this framework. We further evaluated its effectiveness on two widely used PGs — HNSW and Vamana — by creating two variants, HNSW+ and Vamana+, through the integration of our edge pruning rule. Figure 10 presents the recall-NDC trade-offs, with $\alpha$-CNG included for reference. Both HNSW+ and Vamana+ consistently outperformed their original counterparts. HNSW+ achieved more substantial gains over HNSW and even surpassed $\alpha$-CNG on MSONG. We attribute this difference to the distinct pruning strategies used by Vamana and HNSW: Vamana employs a scaled triangle inequality (Fig. 1b), which incorporates the parameter $\alpha$, while HNSW uses the original triangle inequality (Fig. 1a) with a stricter pruning condition. These experimental results demonstrate the transferability of our edge pruning rule to enhance the query performance of other PG solutions.
@@ 6. Related Work
    Approximate nearest neighbor (ANN) search has attracted considerable attention over the past two decades, leading to the development of diverse methodologies aimed at enhancing search performance. Recent experimental studies (abf20; lzah20; lzs+19; wxyw21; aep25) indicate that PG-based approaches surpass other techniques, including hash-based methods (im98; swq+14; hfz+15; wplp24; tzz24; lk20; lwwk20), tree-based methods (ks97; b75; c08; y93; wwj+13; laja08), and inverted index-based methods (bl14; jds10).
@@
    PG-based methods. The Delaunay Graph (DG) (a91), one of the earliest PGs, is the dual graph of the Voronoi diagram with $O(n^{\lceil m/2\rceil})$ space in the $m$-dimensional space. DG guarantees finding the exact NN but suffers from high out-degrees and unbounded search time.
@@
    Several PGs in the literature provide non-trivial query accuracy guarantees while maintaining graph sparsity. Inspired by the relative neighborhood graph (jt92), Arya et al. (am93) and Fu et al. (fxwc19) proposed MRNG. However, it can find the exact NN only when $q\in P$. Fu et al. (fwc22) introduced the satellite system graph (SSG) to support the case when $q\notin P$ and the input is randomly distributed, although the worst-case time remains unbounded. Harwood et al. (hd16) studied the scenario when $\delta(q,v^{*})$ is at most a constant $\tau>0$; when $P$ is uniformly drawn from m, their PG can find the exact NN in $O(n^{2/m}(\ln n)^{2})$ time. Later, $\tau$-MNG (pcc+23) improved the search time to $O(n^{1/m}(\ln n)^{2})$. Recently, Indyk et al. proved that the slow preprocessing version of Vamana guarantees to find an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN of $q$ in $O(\alpha^{d}\cdot\log\Delta\cdot\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ time, regardless of whether $\delta(q,v^{*})\leq\tau$. Refer to other theoretical works studying cases where $q\in P$ (dgm+24) or the data follows specific distributions (l18; ps20).
@@
    The worst-case construction time of all the aforementioned PGs is $\Omega(n^{2})$. To reduce construction time, numerous practical PGs have been proposed in the literature (see (zxc+23; sds+19; fxwc19; my20; hd16; gl23; lkxi21; pcc+23; zth+23; czk+24; wxy+24; ggxl25; yxl+24) and the references therein). A recent survey by Azizi et al. (aep25) identifies several design paradigms for PGs and indicates that neighbor selection (Definition 1) is crucial for improving search performance, warranting further theoretical exploration. This paper aims to contribute to the understanding of this direction.
@@
    Additional directions. Research has explored methods to extend ANN search by incorporating attribute constraints, enabling data retrieval that satisfies both vector similarity and user-specified attributes (gks+23; wlx+23; pkgz24; zqz+24; cscz24; xgg+24). Other studies have focused on the efficient construction of PGs (zth+23; dcl11; om23; yxl+24) and on supporting updates (ssks21; xmb+25; www+20). Recently, quantization-based techniques (jds10; gsl+20; pel+22; abh+23; gl24; ggy+25) have been developed to compress high-dimensional vectors, thereby accelerating distance computations. These techniques can be integrated into PGs to enhance both construction and search efficiency, as investigated in recent research (ggxl25; ssks21).

### v5

@@ 3.3.1. Space and Construction Time
    where the last equality used the assumption $d=O(1)$. Since the minimum distance between any two points in $P$ is $1$, each ball of radius $\frac{1}{3}$ contains at most one point (see Figure 2(b)). Thus, $|\text{Ring}_{i}|=O((\alpha\cdot\tau)^{d})$, meaning that $p$ has $O((\alpha\cdot\tau)^{d})$ out-neighbors in $\text{Ring}_{i}$.
@@
    Similarly, we can obtain $\mathit{\delta}(p,v)>\mathit{\delta}(u,v)\cdot\alpha+(\alpha+1)\tau$. According to the pruning rule defined by (1), at most one of $u$ and $v$ can be in the shortcut set $S$ of $p$. Hence, for any ball $B$ with radius $r_{i}$, at most one point in $\text{Ring}_{i}\cap B$ can be an out-neighbor of $p$. For instance, in Figure 2(c), when $v$ and $v^{\prime}$ are the our-neighbors of $p$, then edges $(p,u)$ and $(p,u^{\prime})$ will not exist in our PG.
@@ 3.3.2. Query Time
    Now, we prove that $\mathit{\delta}(p_{\ell-1},q)\geq 1/2$. If $\mathit{\delta}(v^{*},q)\geq 1/2$, then, since $v^{*}$ is the exact NN of $q$ in $P$, we have $\mathit{\delta}(p_{\ell-1},q)>\mathit{\delta}(q,v^{*})>1/2$. On the other hand, if $\mathit{\delta}(v^{*},q)<1/2$, by triangle inequality and the fact that the minimum pairwise distance in $P$ is $1$, we obtain
@@
    When $\mathit{\delta}(q,v^{*})>\tau$. We say a proximity graph $G^{\prime}$ is $\alpha$-shortcut reachable if for every two vertices $p,z$ of $G^{\prime}$ such that $(p,z)$ is not in $G^{\prime}$, then, there exists an edge $(p,p^{\prime})$ in $G^{\prime}$ such that $\mathit{\delta}(p^{\prime},z)\leq\mathit{\delta}(p,z)/\alpha$. Indyk and Xu (Indyk and Xu, 2023, Theorem 3.4) proved that given an $\alpha$-shortcut reachable PG, a greedy routing starting at any vertex can answer an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN query after visiting $O(\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ hop vertices. Our $\alpha$-CG $G$ is also $\alpha$-shortcut reachable:
@@ 4. Alpha-Convergent Neighborhood Graph with Efficient Construction
    Despite the superior asymptotic query performance of the $\alpha$-CG, its construction time is $\Omega(n^{2})$ in the worst case. In line with existing works such as NSG (Fu et al., 2019), Vamana (Subramanya et al., 2019), and $\tau$-MNG (Peng et al., 2023), we propose a practical variant that approximates $\alpha$-CG. This section also explores strategies for adaptively pruning candidates (setting the parameter $\alpha$) and constructing the index efficiently.
@@ 4.1. The Overall Framework
    We employ a standard framework utilized in existing PG methods, such as NSG (Fu et al., 2019) and $\tau$-MNG (Peng et al., 2023), to generate a local-neighbor set (as opposed to $P\setminus{p}$) as the candidate set for each point $p\in P$. Then, we apply an adaptive pruning strategy (Section 4.2) to select up to $M$ points as the out-neighbors of $p$, where $M$ is a predefined maximum out-degree that ensures efficient graph storage and computation. We refer to resulting graph $G$ as the $\alpha$-convergent neighborhood graph ($\alpha$-CNG). This framework consists of the following phases, as summarized in Algorithm 3.
    The initial phase constructs an approximate $K$-NN graph $G_{0}$, where each data point is connected to its approximate $K$-NNs in $P$. Efficient implementations are available for constructing such graphs, e.g., (Harwood and Drummond, 2016; Dong et al., 2011). We then identify the navigating node $s$ by performing a beam search, starting from a random point with the centroid of $P$ (the geometric mean of all points in $P$) as the query point. The returned NN is designated as $s$.
@@
    Subsequently, backward edges are added to the graph for each edge inserted in the previous phase, ensuring bidirectional connectivity. If any vertex $p$’s out-degree exceeds $M$, its out-neighbor set is pruned accordingly using our adaptive pruning method. Finally, we verify the graph connectivity by performing a depth-first search (DFS) starting from $s$. To ensure that all nodes are reachable from $s$, we add necessary edges for any nodes not included in the DFS tree, in line with the NSG method (Fu et al., 2019).
@@ 4.3. Efficient Graph Construction
    Total construction time. The construction time of $\alpha$-CNG is dominated by two parts: (i) approximate $K$-NN graph construction, and (ii) candidates pruning for all $p\in P$. We utilize the NN-decent algorithm (Dong et al., 2011) to compute the approximate $K$-NN graph, whose empirical time complexity is sub-quadratic.
@@ 4.4. Discussions
    Although $\tau$ is a problem parameter, in our experiments, we set $\tau$ to a small value within the range $[0,10]$ (and in most cases, $\tau$ is smaller than 1) to avoid the large maximum out-degree. Both $\tau$-MNG (Peng et al., 2023) and our paper found that the optimal settings for $\tau$ may vary across different datasets. Hence, we apply a grid search on the test queries. In practical applications, historical queries can be leveraged to determine this parameter. Specifically, we first compare the results obtained with $\tau=0$ against those with $\tau\in\{10,\,1,\,0.1,\,0.01,\,0.001\}$ to identify a coarse range of $\tau$, and then fine-tune $\tau$ to locate the best value.
    Updates. This paper focuses on the development of static PGs that enhance both query accuracy and query time. Supporting updates is beyond the scope of this work. Nonetheless, concepts from existing research—such as periodic global rebuilding (Peng et al., 2023), lazy deletion with masking (Singh et al., 2021), and in-place update strategies (Xu et al., 2025)—could be applied to our $\alpha$-CNG and may be considered for future work.
    5. Experimental Evaluation
@@
    SIFT 128 1M 10K (Jegou et al., 2010) Image
    CRAWL 300 1.98M 10K (Mikolov et al., 2018) Text
    WIKI 384 1M 1K (Merity et al., 2016) Text
    MSONG 420 1M 200 (Bertin-Mahieux et al., 2011) Audio
    LAION-I2I 768 1M 10K (Schuhmann et al., 2022) Image
    GIST 960 1M 1K (Jegou et al., 2010) Image
    DEEP100M 96 100M 10K (Babenko and Lempitsky, 2016) Image
    BIGANN100M 128 100M 10K (Simhadri et al., 2022) Image
    5.1. Experiment Settings
@@
    Datasets. We utilized eight real-world datasets, which are widely adopted in ANN search evaluation (Peng et al., 2023; Gou et al., 2025; Cai et al., 2024; Fu et al., 2021, 2019; Malkov and Yashunin, 2020; Yang et al., 2024). These datasets include six at the 1M scale and two at the 100M scale. The 1M-scale datasets span various application domains, including image (SIFT, LAION-I2I, GIST), audio (MSONG), and text (WIKI, and CRAWL). We also assessed the scalability of our methods using the 100M-scale datasets, including DEEP100M [^5] DEEP100M consists of the first 100 million vectors from the public DEEP1B(Babenko and Lempitsky, 2016) dataset. and BIGANN100M. Table 2 summarizes the key statistics, including the dimensionality (Dim.), the number of base points (# Base), the number of query points (# Query), data source, and data type.
    Competing Methods. Our first method, $\alpha$-CNG, employs an adaptive pruning strategy (Section 4.2). To evaluate the effectiveness of this strategy, we examined Fixed-$\alpha$-CNG, a variant that employs a global $\alpha$ for all data points. Since previous studies (Li et al., 2019; Wang et al., 2021) have consistently shown that PG-based methods outperform non-PG methods such as LSH and IVF, we focus our comparisons on PG-based baselines. We compared our methods against four state-of-the-art PG algorithms: HNSW (Malkov and Yashunin, 2020), Vamana (Subramanya et al., 2019), NSG (Fu et al., 2019), and $\tau$-MNG (Peng et al., 2023), selected for their robust performance. Three additional popular algorithms—NSSG (Fu et al., 2021), DPG (Li et al., 2019), and FANNG (Harwood and Drummond, 2016)—were excluded as $\tau$-MNG (Peng et al., 2023) outperforms their performance. All baseline implementations use publicly available source code.
    The construction parameters for all structures were selected based on official recommendations and empirical evaluations. HNSW used a configuration of $M$ = 32 and $ef_{C}$ = 500 for all datasets, while Vamana used the settings in (Subramanya et al., 2019), with $M$ = 70, $L$ = 75, and $\alpha$ = 1.2.
@@ 5.2. Search Performance
    Recall vs. NDC. We first assessed the trade-off between (average) recall and (average) NDC on the six datasets at the 1M scale. Figure 5 reports the results obtained by varying the queue size $L$ of beam search. Our $\alpha$-CNG outperformed all baselines, achieving higher recall with fewer NDCs. The only exception is the SIFT dataset, for which all methods exhibited strong performance due to its low intrinsic dimensionality (Peng et al., 2023). Fixed-$\alpha$-CNG also surpassed the four baselines, except on MSONG, but was outperformed by $\alpha$-CNG.
@@
    Recall vs. # hops. We next evaluated the (average) number of hops by varying $L$. Figure 5 shows that $\alpha$-CNG substantially reduced the number of hops across all datasets, and Fixed-$\alpha$-CNG also outperformed the baselines on most datasets. Table 3 confirms that the reductions in the number of hops over the best-performing baseline are substantial: exceeding 45% on five out of the six datasets, with a maximum speedup of 2.88x. These gains arise from two key factors: $\alpha$-CNG approximates $\alpha$-CG for faster convergence, and our adaptive pruning strategy preserves more long-distance shortcut edges, effectively reducing the number of hops. The reduced hops of $\alpha$-CNG are advantageous for disk-based (Subramanya et al., 2019) or distributed deployments of PGs, where hops correspond to I/O operations.
@@ 5.3. Effect of Parameters
    Varying $\tau$. Figure 8 illustrates the impact of the parameter $\tau$ on search performance, where the middle tested $\tau$ for each dataset is the optimal $\tau$ identified. We observe a trend similar to that reported for $\tau$-MNG (Peng et al., 2023): as $\tau$ increases from zero, search performance initially improves before deteriorating with further increases in $\tau$. Recall that the search time is proportional to the total out-degrees of the hop vertices. A moderate increase in $\tau$ allows connecting more candidates and enhances the graph connectivity, thus reducing the number of hops. However, if $\tau$ increases excessively, the out-degree may become too large, as the search must compute distances for more neighbors at each step. Additionally, the increased out-degrees may exceed the limit $M$ for some nodes, resulting in the replacement of long-distance shortcut edges with less useful ones. Overall, a small positive $\tau$ enhances graph connectivity and improves search performance, while excessively large values can lead to over-inclusion and reduced pruning effectiveness.
@@ 5.4. Construction Performance
    Table 4. Index construction time and index size across six datasets.
    Index time (secs) Index size (MB)
    HNSW 42 159 91 145 116 206 256 509 256 254 256 256
    Vamana 9 43 19 38 35 67 163 409 114 151 134 118
    NSG 43 174 97 223 93 200 115 183 113 97 83 90
    $\tau$-MNG 43 182 98 242 98 217 113 329 124 117 127 131
    Fixed-$\alpha$-CNG 42 206 88 215 99 207 158 512 179 203 186 162
    $\alpha$-CNG 53 221 119 227 110 228 188 563 228 225 246 227
@@ 5.5. Effectiveness of Our Edge Pruning Rule
    Our two methods, NSG, and $\tau$-MNG, share a framework that generates candidate sets from approximate $K$-NN graphs. Section 5.2 shows the efficacy of our pruning rule (Definition 2) within this framework. We further evaluated its effectiveness on two widely used PGs — HNSW and Vamana — by creating two variants, HNSW+ and Vamana+, through the integration of our edge pruning rule. Figure 9 presents the recall-NDC trade-offs, with $\alpha$-CNG included for reference. Both HNSW+ and Vamana+ consistently outperformed their original counterparts. HNSW+ achieved more substantial gains over HNSW and even surpassed $\alpha$-CNG on MSONG. We attribute this difference to the distinct pruning strategies used by Vamana and HNSW: Vamana employs a more relaxed pruning rule by setting $r=\mathit{\delta}(p,u)/\alpha$ (Fig. 1b), which incorporates the parameter $\alpha$, while HNSW uses the strict pruning rule with $r=\mathit{\delta}(p,u)$ (Fig. 1a). These experimental results demonstrate the transferability of our edge pruning rule to enhance the query performance of other PG solutions.
@@ 6. Related Work
    Approximate nearest neighbor (ANN) search has attracted considerable attention over the past two decades, leading to the development of diverse methodologies aimed at enhancing search performance. Recent experimental studies (Aumüller et al., 2020; Li et al., 2020, 2019; Wang et al., 2021; Azizi et al., 2025) indicate that PG-based approaches surpass other techniques, including hash-based methods (Indyk and Motwani, 1998; Sun et al., 2014; Huang et al., 2015; Wei et al., 2024; Tian et al., 2024; Lu and Kudo, 2020; Lu et al., 2020), tree-based methods (Katayama and Satoh, 1997; Bentley, 1975; Cayton, 2008; Yianilos, 1993; Wang et al., 2013; Lejsek et al., 2008), and inverted index-based methods (Babenko and Lempitsky, 2014; Jegou et al., 2010).
@@
    PG-based methods. The Delaunay Graph (DG) (Aurenhammer, 1991), one of the earliest PGs, is the dual graph of the Voronoi diagram with $O(n^{\lceil m/2\rceil})$ space in the $m$-dimensional space. DG guarantees finding the exact NN but suffers from high out-degrees and unbounded search time.
@@
    Several PGs in the literature provide non-trivial query accuracy guarantees while maintaining graph sparsity. Inspired by the relative neighborhood graph (Jaromczyk and Toussaint, 1992), Arya et al. (Arya and Mount, 1993) and Fu et al. (Fu et al., 2019) proposed MRNG. However, it can find the exact NN only when $q\in P$. Fu et al. (Fu et al., 2022) introduced the satellite system graph (SSG) to support the case when $q\notin P$ and the input is randomly distributed, although the worst-case time remains unbounded. Harwood et al. (Harwood and Drummond, 2016) studied the scenario when $\delta(q,v^{*})$ is at most a constant $\tau>0$; when $P$ is uniformly drawn from m, their PG can find the exact NN in $O(n^{2/m}(\ln n)^{2})$ time. Later, $\tau$-MNG (Peng et al., 2023) improved the search time to $O(n^{1/m}(\ln n)^{2})$. Recently, Indyk et al. proved that the slow preprocessing version of Vamana guarantees to find an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN of $q$ in $O(\alpha^{d}\cdot\log\Delta\cdot\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ time, regardless of whether $\delta(q,v^{*})\leq\tau$. Refer to other theoretical works studying cases where $q\in P$ (Diwan et al., 2024) or the data follows specific distributions (Laarhoven, 2018; Prokhorenkova and Shekhovtsov, 2020).
@@
    The worst-case construction time of all the aforementioned PGs is $\Omega(n^{2})$. To reduce construction time, numerous practical PGs have been proposed in the literature (see (Zhang et al., 2023; Subramanya et al., 2019; Fu et al., 2019; Malkov and Yashunin, 2020; Harwood and Drummond, 2016; Gao and Long, 2023; Lu et al., 2021; Peng et al., 2023; Zhao et al., 2023; Chen et al., 2024; Wang et al., 2024; Gou et al., 2025; Yang et al., 2024) and the references therein). A recent survey by Azizi et al. (Azizi et al., 2025) identifies several design paradigms for PGs and indicates that neighbor selection (Definition 1) is crucial for improving search performance, warranting further theoretical exploration. This paper aims to contribute to the understanding of this direction.
@@
    Additional directions. Research has explored methods to extend ANN search by incorporating attribute constraints, enabling data retrieval that satisfies both vector similarity and user-specified attributes (Gollapudi et al., 2023; Wang et al., 2023; Patel et al., 2024; Zuo et al., 2024; Cai et al., 2024; Xu et al., 2024). Other studies have focused on the efficient construction of PGs (Zhao et al., 2023; Dong et al., 2011; Ono and Matsui, 2023; Yang et al., 2024) and on supporting updates (Singh et al., 2021; Xu et al., 2025; Wei et al., 2020). Recently, quantization-based techniques (Jegou et al., 2010; Guo et al., 2020; Paparrizos et al., 2022; Aguerrebere et al., 2023; Gao and Long, 2024; Gao et al., 2025) have been developed to compress high-dimensional vectors, thereby accelerating distance computations. These techniques can be integrated into PGs to enhance both construction and search efficiency, as investigated in recent research (Gou et al., 2025; Singh et al., 2021).